OpenStudy (anonymous):

Solve the following system of equations. 2x + 2y + z = 10 3x - y + 3z = 10 2x + 3y - 2z = 6 I'm not good at systems of three equations. Anyone willing to help me start?

OpenStudy (anonymous):

From what I attempted: 2x+2y+z=10 3x-y+3x=10 I multiplied the bottom equation by negative three and I got the equation: 8x+7z=30 after adding them. Was I supposed to do this first?

OpenStudy (anonymous):

Can you tell me what you did to get those two equations?

OpenStudy (anonymous):

Also, with those two equations do I use them together to solve for one of the variables?

OpenStudy (anonymous):

2x + 2y + z = 10 3x - y + 3z = 10 -->(2) -------------- 2x + 2y + z = 10 6x - 2y + 6z = 20 (result of multiplying by 2) --------------add 8x + 7z = 30 3x - y + 3z = 10 -->(3) 2x + 3y -2z = 6 ------------- 9x - 3y + 9z = 30 (result of multiplying by 3) 2x + 3y - 2z = 6 ------------add 11x + 7z = 36 8x + 7z = 30 --->(-1) 11x + 7z = 36 ------------ -8x -7z = -30 (result of multiplying by -1) 11x + 7z = 36 -----------add 3x = 6 x = 2 8x + 7z = 30 8(2) + 7z = 30 16 + 7z = 30 7z = 30 - 16 7z = 14 z = 2 2x + 2y + z = 10 2(2) + 2y + 2 = 10 4 + 2y + 2 = 10 2y + 6 = 10 2y = 10 - 6 2y = 4 y = 2 check... 3x - y + 3z = 10 3(2) - 2 + 3(2) = 10 6 - 2 + 6 = 10 12 - 2 = 10 10 = 10 (correct) x,y,and z, are all 2

OpenStudy (anonymous):

Thanks to both of you! I'll use Kellie's work for reference on any more problems, and thanks DSS for starting me off

OpenStudy (anonymous):

you have to be really careful with elimination because if you mess up at the beginning, that mistake carries all the way through to the end

OpenStudy (anonymous):

Thanks for the advice, it'll help a lot since Elimination is the main form I use to solve these

OpenStudy (anonymous):

just remember, whatever variable you eliminate in the first two equation you use, you have to eliminate the same variable in the second set of equations you use.

OpenStudy (anonymous):

I would suggest checking your answers at the end to make sure they are correct

OpenStudy (anonymous):

okay. By the way, do you know a link to any good Logarithm videos?

OpenStudy (anonymous):

no sorry...not really

OpenStudy (anonymous):

good luck finding some :) and I am glad I could help with the elimination